# A lower bound on the number of bent squares

### v2

@@ 1 Introduction
$\operatorname{log}_{2}b_{n}\geq cn\cdot 2^{\frac{n}{2}}(1+o(1))$
with $c=\frac{1}{2}$, and showed that a construction by Baksova and Tarannikov [2] yields $c=\frac{3}{4}$.
Agievich [3] introduced the representation of a bent function as a bent square. In this note, we apply this representation to obtain $c=1$.
@@ 2 Preliminaries
A bent function in $n$ variables is a Boolean function $f$ in $n$ variables such that the Walsh transform
@@ 3 Bent squares
The number of $2^{\frac{n}{2}}\times 2^{\frac{n}{2}}$ bent squares whose rows and columns consist of vectors of type 1 is easily seen to be $2^{2^{\frac{n}{2}}}\left(2^{\frac{n}{2}}\right)!$. We deviate from [3] in the estimate of the number of $2^{\frac{n}{2}}\times 2^{\frac{n}{2}}$ bent squares whose rows and columns consist of vectors of type 2.
@@ 4 A lower bound on $b_{n}$
For each of these matrices, there are at least 32 ways to replace the non-zero entries by $\pm 2^{\frac{n}{2}}$ so that we end up with an odd number of positive entries in each row and each column. For example, we can let the entries of AA be all positive or all negative, let $B$ and $C$ have one positive and one negative entry in each row and each column, and let the entries of $D$ be all positive or all negative. This accounts for at least 16 ways, by Lemma 4.2. In addition, we can let AA and $D$ have one positive and one negative entry in each row and each column, while the entries of $B$ and of $C$ are either all positive or all negative.

### v3

@@ 1 Introduction
$\operatorname{log}_{2}b_{n}\geq cn\cdot 2^{\frac{n}{2}}(1+o(1))$ (1)
with $c=\frac{1}{2}$, and showed that a construction by Baksova and Tarannikov [2] yields $c=\frac{3}{4}$. The currently best upper bound is
$\operatorname{log}_{2}b_{n}\leq\frac{3}{8}\cdot 2^{n}(1+o(1))$
by Potapov [4].
Agievich [1] introduced the representation of a bent function as a bent square, that is, a square matrix for which each row and each column is the Walsh spectrum of some Boolean function (confer Section 3). In Section 4, we apply this representation in order to prove that (1) holds with $c=1$. The result is not based on an explicit construction. Rather, we consider a set of potential building blocks, and show that a sufficiently large number of combinations must yield valid bent squares.
@@ 2 Preliminaries
A Boolean function $f$ in $n$ variables is called bent if the Walsh transform
@@ 3 Bent squares
The number of $2^{\frac{n}{2}}\times 2^{\frac{n}{2}}$ bent squares whose rows and columns consist of vectors of type 1 is easily seen to be $2^{2^{\frac{n}{2}}}\left(2^{\frac{n}{2}}\right)!$, which implies (1) with $c=\frac{1}{2}$. We deviate from [1] in the estimate of the number of $2^{\frac{n}{2}}\times 2^{\frac{n}{2}}$ bent squares whose rows and columns consist of vectors of type 2.
@@ 4 A lower bound on $b_{n}$
For each of these matrices, there are at least 32 ways to replace the non-zero entries by $\pm 2^{\frac{n}{2}-1}$ so that we end up with an odd number of positive entries in each row and each column. For example, we can let the entries of AA be all positive or all negative, let $B$ and $C$ have one positive and one negative entry in each row and each column, and let the entries of $D$ be all positive or all negative. This accounts for at least 16 ways, by Lemma 4.2. In addition, we can let AA and $D$ have one positive and one negative entry in each row and each column, while the entries of $B$ and of $C$ are either all positive or all negative.